Rose is 25 years younger than her father. Five years ago, her father was twice as old as she will be 5 years from now. How old is Rose today?
Answer by lwsshak3(11628)

Rose is 25 years younger than her father. Five years ago, her father was twice as old as she will be 5 years from now. How old is Rose today
**
let x=Rose's age today
x+25=Father's age today
5 years ago:
Father's age=x+25-5=x+20
5 years from now:
Rose's age=x+5
..
(x+20)/(x+5)=2
2x+10=x+20
2x-x=20-10
x=10
ans:
Rose's age today=10